### v2

@@ 3.1 System Architecture
Deterministic ML Execution. To combat the inherent inconsistencies caused by floating-point calculations in ML, our network implements fixed-point arithmetic and software-based floating-point libraries. To fix the randomness, both the asserter and the validator will be assigned the same random seed. This approach ensures uniform, deterministic ML executions, enabling the use of a deterministic state transition function for the ML process, enhancing reliability in decentralized environments.
Stateless Design in ML Inference. While applications such as inference might appear stateful to users—due to the ability to engage in ongoing interactions—they are fundamentally stateless. In our framework, each query is treated as independent; any necessary historical context is encapsulated within each new request, thereby maintaining statelessness throughout the ML process.
Permissionless Network Participation. In our network, anyone can join the network, gain full access to the ML model $f(x)$, and contribute by running an AI server. The function $f(x)$, pivotal for the AI inference process, is established as common knowledge within the network. This inclusivity ensures that the model is capable for validation, promoting the security of the network.
@@
Off-chain Operations. Given the extensive usage inherent in AI inference networks, it is impractical to process or record every AI inference outcome on the blockchain due to scalability constraints. Instead, AI inferences are computed off-chain by decentralized servers, which then relay the results along with their digital signatures directly to users, bypassing the on-chain mechanism. This approach significantly reduces the load on the blockchain while ensuring that users receive authenticated and accurate inference results.
@@
On-chain Operations. Critical functions, such as the posting of overall balance calculations after set periods, are conducted on-chain to ensure transparency and security. Additionally, the protocol allows for on-chain handling of challenge mechanism, enabling disputes or anomalies detected off-chain to be resolved transparently and securely within the blockchain framework. This ensures both integrity and accountability in the system’s operations.
3.2 SpML Protocol
The spML protocol is designed as follows.
@@ 3.2 SpML Protocol
    The user sends the input to a randomly selected server A with his signature. The base fee is $u$, which is determined by the protocol, and will be burnt.
@@
    Server A sends the output and its hash back to the user with his signature.
@@
    With a predetermined probability $p$, the protocol randomly select another server B. Otherwise, A gets the reward $R$, and this round of AI inference concludes.
@@
    If selected, user sends the same input to server B with his signature.
@@
    Server B sends the output and its hash back to the user with his signature.
@@
    The user compares the hash of A’s and B’s output. If they are the same, user accepts the output and gets a pre-determined discount $d<1$ of his base fee, which equals to $d\cdot u$, back to his account, both of A and B gets the reward $R$, and this round of AI inference concludes.
@@
    Otherwise, the user broadcast A’s and B’s hash to the network. Server A and B also broadcast the user’s input to the network, to prevent the user sends different input to them.
@@
    All the nodes vote for the behavior of user, sever A and server B. A result will be considered as valid if at least $2/3$ of the nodes vote on it. If the user is dishonest, the user will be slashed $S_{u}$. Otherwise, the user will get the transaction fee and the compensation $u+S/3$, the dishonest server will get slashed $S$ and the honest server will get rewarded $R$.
The user and the voting procedure effectively acts as the orchestrator and arbitration procedure respectively, as in PoSP in Section 2.1, due to the comprehensive participation of network nodes. The integrity of the voting outcome is upheld with the same level of trust as that observed in Ethereum Layer 1. Unlike conventional approaches that may require all nodes to recompute the function $f(x)$ during fraud detection processes, spML protocol simplifies this by necessitating only hash comparisons in most cases. This task parallels the execution of a rudimentary smart contract, incurring little computational overhead. Consequently, our network does not experience significant additional load during challenge mechanism operations.
@@ 3.3 Analysis
Property 2
The rational user will never collude with the server in the network, and the dominant strategy for the user is always act honestly.
This is because it is always impossible to slash the honest node. Then we can assume that the user will not collude with any server in the network. Moreover, if the challenge mechanism is triggered, the rational user will indeed send his input to another server B, because he has the utility to gain the transaction fee discount and get to know the true outcome of his request.
Given Property 2, we can assume that the user will always act honestly. In this case, the orchestrator, i.e. the user, is honest, and the spML protocol in Section 3.2 will reduce to the PoSP protocol in Section 2.1.
@@
$p>\frac{C}{(1-r)(R+S)},$
the system has a unique Nash Equilibrium in pure strategies, where every participant perform honestly.
@@
After plugging $n=1$, $R_{A}=R_{V}=R$, $U_{1}=R$ and $U_{2}=2R$ in Theorem 2.1, we can get this result.
@@
zkML. At its core, zkML leverages zero-knowledge proofs. In the context of decentralized AI, zkML ensures that computations can be verified for correctness without revealing the underlying data or the specifics of the computation. This characteristic is particularly advantageous for applications requiring stringent data privacy measures. However, the sophistication and computational intensity of generating zero-knowledge proofs present challenges in terms of scalability and accessibility.
@@
Security More vulnerable to fraudulent activity High security through cryptographic proofs High security through economic incentives
@@
Delays Potential delays in dispute resolution Delays due to proof generation Almost no delay due to asserter will act honestly
@@
Scalability Highly scalable Limited by computational overhead of proof generation Highly scalable
Simplicity Simple, unless for fraud proof Complex due to zk proofs Simple
@@
Overhead Low computational overhead, unless in the case of disputes High computational overhead due to the nature of cryptographic proof generation Low computational overhead, unless in the case of disputes which never happens if everyone is rational
The table above compares opML, zkML and spML as follows.
Security. OpML is recognized as being more vulnerable to fraudulent activity, a vulnerability that can arise from insufficient validation of results. Typically, users themselves may initiate a fraud proof if the results they receive appear aberrant. Nevertheless, it presents a challenge for users to discern the accuracy of these results, particularly when a server might employ a similar yet less sophisticated model to produce the output. Conversely, zkML boasts robust security due to its use of cryptographic proofs. SpML can also achieve high-security just like zkML, but achieves this through economic incentives. In spML, the initiation of challenge mechanism is an automated process managed by the protocol itself, rather than relying on the assumption that there will be at least one external validator, as is the case with opML. This intrinsic mechanism enhances the overall security of our system.
@@
Delays. In opML, delays exist due to the challenge period, during which a transaction can be challenged with a fault proof. This is a drawback in scenarios requiring real-time results. zkML faces inherent significant delays due to the computational overhead in proof generation. SpML is designed to mitigate delay issues altogether. Even if the challenge mechanism is triggered, the user does not need to wait for the challenge procedure, because the result from the asserter is trustable due to the unique Nash Equilibrium in pure strategies.
@@
Scalability. OpML is recognized for its scalability, especially when disputes are minimal, suggesting a lightweight protocol suitable for large-scale applications. ZkML’s scalability is hampered by the heavy computational load required for proof generation. In contrast, spML is presented as highly scalable as well, which can handle extensive network activity without significant degradation in performance.
Simplicity. The complexity of ZK proofs contributes to the high complexity of zkML, possibly making it less accessible for broader implementation. For opML, it is usually considered to be simple. However, its implementation can become complex when applied to fraud proof scenarios. In contrast, spML maintains consistent simplicity in implementation due to its arbitration procedure based on voting. This simplicity enhances ease of integration and maintenance, potentially facilitating widespread adoption.
@@
Overhead. OpML claims low computational overhead, with the caveat that opML may incur higher overhead during disputes. ZkML’s approach results in high computational overhead due to cryptographic processes. SpML also has a low computational overhead. During the challenge mechanism, spML still has a low computational overhead. This is because in spML, the validators only need to compare some hashes in most cases. Only when the hashes do not match during the challenge mechanism, spML may incur high overhead during arbitration, but the arbitration never happens if every node is rational.

### v3

@@ 3.1 Design Principles & Assumptions
To combat the inherent inconsistencies caused by floating-point calculations in ML, we follow similar practices from prior work [40, 7], which implements fixed-point arithmetic and software-based floating-point libraries to ensure determinsitic ML execution. To fix the randomness, both the asserter and the validator will be assigned the same random seed. This approach ensures uniform, deterministic ML executions, enabling the use of a deterministic state transition function for the ML process, enhancing reliability in decentralized environments.
@@
Minimizing On-chain Operations. Given the extensive usage in AI inference networks, processing or recording every AI inference outcome on the blockchain is impractical due to scalability constraints. Instead, AI inferences are computed off-chain by decentralized servers, which then relay the results and their digital signatures directly to users, bypassing the on-chain mechanism. On-chain operations are only necessary during arbitration, which can be avoided if all nodes act economically rational. This approach significantly reduces the blockchain’s load while ensuring users receive authenticated and accurate inference results.
@@
Critical functions, such as posting overall balance calculations at set intervals, are conducted on-chain to ensure transparency and security. Additionally, the protocol allows for on-chain handling of challenge mechanisms, enabling transparent and secure resolution of disputes or anomalies detected off-chain within the blockchain framework. This ensures both integrity and accountability in the system’s operations.
3.2 System Setup and Threat Model
Our system consists of users, executors, orchestrators and a blockchain. An executor chosen to run the query of a user is called an asserter and the one chosen to rerun the query of a user is called a validator.
We assume there are a total of $N$ executor nodes. Each executor is required to deposit at least $S$ in order to be considered as a valid executor. Out of these, we assume that at most a fraction $r$ of them are malicious. A malicious executor can behave in arbitrary ways including colluding with other malicious executors.
We assume there are a total of $13f+$ 1 orchestrators and at most $f$ of them are malicious. The rest of them are trusted for availability and security. The orchestrators implement a Byzantine Fault Tolerance (BFT) consensus with state machine replication to tolerate the Byzantine behavior of the malicious nodes. In our protocol, we assume all actions taken from the orchestrator nodes are under BFT consensus.
In addition, these orchestrator nodes jointly run a distributed random beacon protocol [6] that tolerates up to $f$ malicious nodes to generate a public random value at the beginning of each epoch. Each node can query from this randomness beacon at epoch $t$ and obtain $\tau_{t}$ as the random seed. We assume that the random value is a block cipher key (and if it is not, one can apply a key derivation function).
We further assume that the blockchain is trusted for availability and security. It maintains a trusted list of all executors, a trusted PKI that stores public keys for all users and executors and the balances of all parties. At the beginning of the protocol, users and executor nodes query the trusted PKI to obtain copies of all orchestrator nodes’ public keys. We assume that executors and users have the correct list of public keys for the orchestrators and we discussed in the next version of this document how to perform membership changes. The orchestrator nodes also keep copies of the public keys of all registered users and executor nodes.
Communication between any two entities in this system happens over TLS.
3.3 Notations and Cryptographic Building Blocks.
We use PRF to denote a pseudorandom function [11], which is a function that can be used to generate output from a random seed and a data variable, such that the output is computationally indistinguishable from truly random output. We define two deterministic sampling functions, $\texttt{Bucket}:\mathcal{S}\times\mathbb{N}\to{1,2,\dots,N}$ and $\texttt{Random}:\mathcal{S}\to[0,1)$, as follows:
  • $\texttt{Bucket}(\mathsf{seed},\mathsf{string},N)=\mathsf{PRF}_{\mathsf{seed}}(\mathsf{string})\mod N$: maps a random block cipher key $\mathsf{seed}$ and a unique string to a random but deterministic value in $\{1,\dots,N\}$
  • $\texttt{Sampled}(\mathsf{seed},\mathsf{uniquestr},p)=[\mathsf{PRF}_{\mathsf{seed}}(\mathsf{uniquestr})<p\times\mathsf{PRFMax}$]: returns true with approximately probability $p$ based on a random block cipher key $\mathsf{seed}$ and a unique string.
To denote a digital signature over the message $x$, we use $\sigma_{x}$. To prevent replay attacks, each user request is assigned a unique request ID (reqid). We assume that every digital signature $\sigma_{x}$ is generated over the combination of the original message $x$ and this reqid.
3.4 The SpML protocol
In this section, we describe spML, a concrete instantiation of the PoSP protocol to enable verifiable machine learning inference.
The orchestrators have a record of all status of pending transactions (the transactions that haven’t been finalized and posted to the blockchain yet) and the updates of the balances of the users and the executors.
The user has an input $x$ and wants to compute $f(x)$, where $f$ is a machine learning model. The spML protocol is designed as follows.
Basic Protocol
Step 1: User Request Submission
[…]
@@ 3.4 The SpML protocol
[…]
@@
[…]
@@
    Orchestrators engage in a BFT agreement to accept this reqid. This agreement ensures that the reqid has not been processed before and prevents duplicate processing. If the BFT agreement is successful and the signature is valid, the request $(x,\textsf{reqid})$ is accepted and stored in the pending transaction pool of each orchestrator node.
Step 2: Orchestrator Agreement and Asserter Selection
The orchestrators agree (through BFT) on processing the user’s request $(x,\textsf{reqid})$ at the current request processing epoch $t_{\textsf{req}}$. Each of them does the following:
  1. 1.
    Queries the random seed $\tau_{t_{\textsf{req}}}$ based on this epoch and reuses the stored $\mathsf{pk}_{\textsf{user}}$ from Step 1.
  2. 2.
    Calculates $i:=\texttt{Bucket}(\tau_{t_{\textsf{req}}},\mathsf{pk}_{\textsf{user}}||x||\textsf{reqid},N)$ to select the asserter (node $i$).
  3. 3.
    Signs over $(x,\textsf{reqid})$ (denote orchestrator $k$’s signature as $\sigma_{(x,\textsf{reqid})}^{k}$) and sends $(x,\textsf{reqid},\sigma_{(x,\textsf{reqid})}^{k})$ to the selected asserter, node $i$.
Step 3: Asserter Execution and Response
The asserter, node $i$, upon receiving messages from orchestrators, does the following (within a defined timeout $T_{assert}$):
  1. 1.
    Verifies $\sigma_{(x,\textsf{reqid})}^{k}$ is a valid signature over $(x,\textsf{reqid})$ signed by the orchestrator $k$.
  2. 2.
    If the asserter receives at least $12f+$ 1 such verified messages for the same $(x,\textsf{reqid})$ from distinct orchestrators, it accepts the request and calculates $y_{i}=f(x)$.
  3. 3.
    Signs the result: $\sigma_{i}:=\textsf{Sign}_{\textsf{asserter}_{i}}(x,\textsf{reqid},y_{i})$.
@@
    Sends $(x,\textsf{reqid},y_{i},\sigma_{i})$ back to all orchestrator nodes.
@@
    Handling Asserter Failure: If node $i$ fails to respond to orchestrators with a valid $(x,\textsf{reqid},y_{i},\sigma_{i})$ within timeout $T_{assert}$, orchestrators will BFT-agree on this failure. The request will be re-assigned and the asserter may be penalized (see Timeout and Failure Handling Protocol).
Each orchestrator node, upon receiving $(x,\textsf{reqid},y_{i},\sigma_{i})$ from asserter $i$:
  1. 1.
    Verifies the asserter’s signature $\sigma_{i}$ on $(x,\textsf{reqid},y_{i})$.
  2. 2.
    If valid, forwards $(y_{i},\sigma_{i},\textsf{reqid})$ to the user.
The user can proceed to consume the result $y_{i}$ once they receive $(y_{i},\sigma_{i},\textsf{reqid})$ from at least $12f+$ 1 orchestrators, all containing the same $y_{i}$ value for the given reqid, and after personally verifying at least one valid $\sigma_{i}$.
Step 4: Orchestrator Post-Execution and Challenge Decision
The orchestrators verify the asserter’s signature $\sigma_{i}$ (if not already done). They then agree (through BFT) on processing the asserter’s result $(y_{i},\sigma_{i})$ for request reqid. This occurs before or at the start of the challenge decision epoch $t_{\textsf{chal}}$. Each orchestrator then does the following:
  1. 1.
    Queries the random seed $\tau_{t_{\textsf{chal}}}$ based on this epoch $t_{\textsf{chal}}$ and $\mathsf{pk}_{\textsf{user}}$ from the PKI.
  2. 2.
    Calculates $b:=\texttt{Sampled}(\tau_{t_{\textsf{chal}}},\mathsf{pk}_{\textsf{user}}||x||\textsf{reqid},p)$, where $p$ is a pre-determined threshold.
  3. 3.
    If $b=1$, then it initiates the Challenge Protocol for $(x,\textsf{reqid},y_{i},\sigma_{i})$. Otherwise, it records a reward of $R<B/2$ for the asserter $i$ locally (associated with reqid), and concludes this request’s active processing.
Challenge Protocol
Step 1: Initiation and Validator Selection
If the Challenge Protocol is initiated for $(x,\textsf{reqid},y_{i},\sigma_{i})$, each orchestrator does the following:
  1. 1.
    Uses the same random seed $\tau_{t_{\textsf{chal}}}$ (from Basic Protocol Step 4.1).
  2. 2.
    Calculates $j:=\texttt{Bucket}(\tau_{t_{\textsf{chal}}},\mathsf{pk}_{\textsf{user}}||x||\textsf{reqid},N)$ to select the validator (node $j$). (Note: Asserter $i$ cannot be selected as validator $j$ for the same request. If $j=i$, a deterministic rule, e.g., $j=(i+1)\pmod{N}$, should apply.)
  3. 3.
    Sends $(x,\textsf{reqid})$ along with its orchestrator signature $\sigma_{(x,\textsf{reqid})}^{k}$ (from Basic Protocol Step 2.3) to the validator, node $j$.
Step 2: Validator Execution and Response
The validator, node $j$, upon receiving messages from orchestrators, does the following (within a defined timeout $T_{validate}$):
  1. 1.
    Verifies $\sigma_{(x,\textsf{reqid})}^{k}$ is a valid signature over $(x,\textsf{reqid})$ signed by the orchestrator $k$.
  2. 2.
    If the validator receives at least $12f+$ 1 such verified messages for the same $(x,\textsf{reqid})$ from distinct orchestrators, it accepts the task.
  3. 3.
    Calculates $y_{j}=f(x)$.
  4. 4.
    Signs the result: $\sigma_{j}:=\textsf{Sign}_{\textsf{validator}_{j}}(x,\textsf{reqid},y_{j})$.
  5. 5.
    Sends $(x,\textsf{reqid},y_{j},\sigma_{j})$ back to all orchestrator nodes.
@@
    Handling Validator Failure: If node $j$ fails to respond to orchestrators with a valid $(x,\textsf{reqid},y_{j},\sigma_{j})$ within timeout $T_{validate}$, orchestrators will BFT-agree on this failure. The request will be reassigned and Node $j$ may be penalized. (see Timeout and Failure Handling Protocol).
Step 3: Orchestrator Verification and Outcome
Each orchestrator, upon receiving $(x,\textsf{reqid},y_{j},\sigma_{j})$ from validator $j$:
  1. 1.
    Verifies the validator’s signature $\sigma_{j}$ on $(x,\textsf{reqid},y_{j})$.
  2. 2.
    If valid, it compares $y_{j}$ with the previously received $y_{i}$ for the same reqid.
  3. 3.
    If $y_{i}=y_{j}$: Records a reward of $R<B/2$ for each of the asserter $i$ and the validator $j$ locally (total $2R<B$), and concludes this request’s active processing.
  4. 4.
    Otherwise (if $y_{i}\neq y_{j}$): Initiates the Arbitration Protocol, providing $(x,\textsf{reqid},y_{i},\sigma_{i},y_{j},\sigma_{j})$.
Arbitration Protocol
Step 1: Orchestrator Request to Smart Contract
If the Arbitration Protocol is initiated, each of the orchestrators sends an arbitration request containing $(\texttt{Arbitration},x,\textsf{reqid},y_{i},\sigma_{i},y_{j},\sigma_{j})$ along with its own signature on this entire tuple, to the arbitration smart contract on the blockchain.
Step 2: Smart Contract Verification and Judgment
The smart contract does the following:
  1. 1.
    Queries the public keys of the orchestrators and verifies the signatures on the arbitration requests.
  2. 2.
    If at least $12f+$ 1 valid and identical arbitration requests are received from distinct orchestrators, it continues. Otherwise, it aborts or awaits more identical requests.
  3. 3.
    Verifies that $\sigma_{i}$ is a valid signature of asserter $i$ on $(x,\textsf{reqid},y_{i})$.
  4. 4.
    Verifies that $\sigma_{j}$ is a valid signature of validator $j$ on $(x,\textsf{reqid},y_{j})$.
  5. 5.
    It computes (or securely verifies via ZKP) the true result $y_{true}=f(x)$.
  6. 6.
    Identifies honest and dishonest parties:
    • If $y_{i}=y_{true}$ and $y_{j}\neq y_{true}$: Asserter $i$ is honest, validator $j$ is dishonest.
    • If $y_{i}\neq y_{true}$ and $y_{j}=y_{true}$: Asserter $i$ is dishonest, validator $j$ is honest.
    • If $y_{i}\neq y_{true}$ and $y_{j}\neq y_{true}$ (and $y_{i}\neq y_{j}$): Both are dishonest. (If $y_{i}=y_{j}\neq y_{true}$, challenge should not have led to arbitration, but if it does, both made the same error).
@@
    Slashes the node(s) that generated the wrong result(s) with a certain amount $S$ of their deposit.
@@
    Records the arbitration outcome (who was correct/incorrect, slashed amounts).
Step 3: Orchestrator Post-Arbitration Processing
Each orchestrator queries the arbitration result from the smart contract for reqid.
  • If asserter $i$ was deemed honest, records a reward (e.g., $R_{base}+S_{collected\_from\_j}$) for asserter $i$.
  • If validator $j$ was deemed honest, records a reward (e.g., $R_{base}+S_{collected\_from\_i}$) for validator $j$.
  • If both were dishonest, their deposits $S$ are slashed.
The user is also informed of the arbitrated correct result $y_{true}$.
Timeout and Failure Handling Protocol
Case 1: Asserter Fails to Respond (Basic Protocol Step 3)
If asserter $i$ fails to provide a valid signed result within $T_{assert}$:
  1. 1.
    Orchestrators BFT-agree on the timeout.
  2. 2.
    Reassign asserter: Orchestrators may select a new asserter $i^{\prime}$ (e.g., $i^{\prime}:=\texttt{Bucket}(\tau_{t_{\textsf{req}}},\mathsf{pk}_{\textsf{user}}||x||\textsf{reqid}||\textsf{attempt\_2},N)$). The protocol restarts for this request from Basic Protocol Step 2.3 with the new asserter. Node $i$ may be penalized (e.g., small penalty recorded locally, or via smart contract if persistent).
Case 2: Validator Fails to Respond (Challenge Protocol Step 2)
If validator $j$ fails to provide a valid signed result within $T_{validate}$:
  1. 1.
    Orchestrators BFT-agree on the timeout.
  2. 2.
    Reassign Validator: Similar to asserter failure, the protocol reassigns another validator to complete this request, and validator $j$ may be penalized.
Settlement Protocol
This protocol is executed periodically, e.g., at the end of each settlement epoch $t_{\textsf{settle}}$.
Step 1: Orchestrator Agreement on Balance Updates
After each settlement epoch, each orchestrator:
  1. 1.
    Identifies all transactions (identified by reqid) that have concluded (either via Basic, Challenge, or Arbitration outcomes) within this epoch and whose financial implications (rewards, slashes) have not yet been posted on the blockchain.
  2. 2.
    Calculates the net updates to the balances of all involved users and nodes (asserters, validators) based on recorded rewards and slashes.
  3. 3.
    Crucially, all orchestrators engage in a BFT agreement protocol to arrive at a single, consistent list of balance updates for the epoch.
Step 2: Submission to Smart Contract
Once BFT agreement on the balance updates is reached:
  1. 1.
    One or more (or all) orchestrators submit this agreed-upon batch of balance updates, along with evidence of the BFT agreement (e.g., $12f+$ 1 signatures from orchestrators on the hash of the batch), to the settlement smart contract on the blockchain.
  2. 2.
    The smart contract verifies the BFT agreement proof (e.g., the $12f+$ 1 orchestrator signatures).
  3. 3.
    If verified, the smart contract applies the batch of balance updates to the users’ and nodes’ accounts managed by it.
3.5 Discussion
Malicious Users
Malicious users could collude with malicious executors. If the selected asserter is malicious, then a malicious user could collude with the asserter so that the asserter gains rewards for computing nothing $1-p$ percent of the time. However, the user has to pay for the query and our protocol guarantees that the net reward for the malicious user and asserter is negative.
Malicious Executors
Assuming that a fraction of $r$ executor nodes are malicious, then the probability of selecting two malicious executors is $r^{2}$. In theory, our user could receive two incorrect but same results from the asserter and the validator if they are both malicious.
However, when the asserter responds to the orchestrator, it has no knowledge of and cannot control which node will be selected next. Therefore, if the asserter wants to cheat and provide a wrong response, it can only pass the check $(1-p)+pr$ percent of the time (either the challenge does not happen, or the challenge protocol chooses another malicious node).
Even though malicious executors could choose not to follow the protocol, our mechanism design guarantees that the net reward for them is negative, and any economically rational executor would behave honestly.
Malicious Orchestrators
In our protocol, up to $f$ orchestrator nodes could be malicious. These malicious orchestrators cannot interfere with the correctness of the BFT concensus since there are $12f+$ 1 honest nodes. They could, however, collude with malicious users and executors. If a malicious validator is selected, the malicious orchestrator node could tell the validator the computed f(x) from the asserter. Then, the malicious validator could free-ride the asserter’s result and save the computational cost of computing $f(x)$. Our protocol does not prevent against this cost-saving behavior. However, this behavior does not impact the security of our protocol, since our mechanism design guarantees that any economically rational asserter will compute the correct result. The only case where the user fails to obtain the correct result is when both the asserter and validator are malicious, and the asserter computes the wrong result. As we analyze above, the expected reward for this asserter is negative in this case.
Unresponsive Executors
In our protocol, we assume all executor nodes are available up to a certain limit. An executor node can reject a request if the number of requests it is concurrently handling exceeds this limit, but it is expected to respond to user requests if the limit is not met. If an asserter or validator does not respond within a specified number of epochs, the orchestrators generate a new random seed from the random beacon and send the requests to a new asserter or validator, potentially penalizing the unresponsive node. This setup incentivizes executor nodes to respond promptly. Even if an executor does not respond, the only negative effect for the user is a certain period of delay.
Multi-tier Executors
For a network of nodes with heterogeneous compute power, they can be grouped into tiers of executors based on compute power. This ensures balanced workloads and fair resource allocation, with consistent performance within each tier. Users can choose different tiers based on their requirements and budget, optimizing for either high computational power or cost-efficiency. In the spML protocol, orchestrators assign tasks within the same tier to make sure both the asserter and the validator are able to compute the user’s request $f(x)$, hence enhancing verification robustness and network integrity.
Aggregating Signatures for Reduced Cost
The current protocol design requires an orchestrator to send multiple signatures to the blockchain. It can be reduced to only one signature by using a threshold signature scheme.
@@ 3.6 Analysis
$p>\frac{C}{(1-r)S+(1-2r)R},$
the system has a unique Nash Equilibrium in pure strategies, where every participant outputs the correct result.
@@
After plugging $n=1$, $R_{A}=R$, $U_{1}=R$ and $U_{2}=2R$ in Theorem 2.1, we can get this result.
@@
ZkML. At its core, zkML leverages zero-knowledge proofs. In the context of decentralized AI, zkML ensures that computations can be verified for correctness without revealing the underlying data or the specifics of the computation. This characteristic is particularly advantageous for applications requiring stringent data privacy measures. However, the sophistication and computational intensity of generating zero-knowledge proofs present challenges in terms of efficiency and accessibility.
@@
Security Relies on AnyTrust assumption (no pure-strategy Nash equilibrium) High security through cryptographic proofs Security through economic incentives (pure-strategy Nash equilibrium)
@@
Delays Potential delays in dispute resolution Delays due to proof generation Almost no delay since rational asserter will act honestly
@@
Efficiency Highly efficient Limited by computational overhead of proof generation Highly efficient
@@
Overhead Low computational overhead, unless in the case of disputes High computational overhead due to the nature of cryptographic proof generation Low computational overhead, unless in the case of disputes which never happen if everyone is rational
Table 2: Comparison of OpML, ZkML and SpML
Table 2 compares opML, zkML and spML.
Security. The security of opML relies on the AnyTrust assumption. OpML only has a mixed strategy Nash Equilibrium, which means that there is a positive probability for undetected fraud if every node is rational. Conversely, zkML boasts robust security due to its use of cryptographic proofs. The security of spML is based on economic incentives. In spML, the initiation of challenge mechanism is an automated process managed by the protocol itself, rather than relying on the assumption that there will be at least one external validator, as is the case with opML. The pure strategy Nash Equilibrium demonstrated by spML offers evidence that the system can be deemed secure, provided that each node behaves rationally.
@@
Delays. In opML, delays exist due to the challenge period, during which a transaction can be challenged with a fraud proof. This is a drawback in scenarios requiring real-time results. zkML faces inherent significant delays due to the computational overhead in proof generation. SpML is designed to mitigate delay issues altogether. Even if the challenge mechanism is triggered, the user does not need to wait for the challenge procedure: the user can trust the result because the dominant strategy for the asserter is to output the correct result, if every node is rational.
@@
Efficiency. OpML is recognized for its efficiency, especially when disputes are minimal, suggesting a lightweight protocol suitable for extensive applications. ZkML’s efficiency is hampered by the heavy computational load required for proof generation. In contrast, spML is presented as highly efficient as well, which can handle extensive network activity without significant degradation in performance.
@@
Overhead. OpML claims low computational overhead, with the caveat that opML may incur higher overhead during disputes. ZkML’s approach results in high computational overhead due to cryptographic processes. SpML also has a low computational overhead. During the challenge mechanism, spML still has a low computational overhead. This is because in spML, the challenge mechanism happens very rarely. Only when the results do not match during the challenge mechanism, spML may incur high overhead during arbitration, but the arbitration never happens if every node is rational.